# Who’s asking? User personas and the mechanics of latent misalignment

### v1

@@ H.1 Procedure for generating persona data
We follow a procedure similar to the one described in §G. We used text-unicorn@001 via Cloud Vertex AI to create 100 first-person statements of identity that conform with a particular persona. Specifically, we prompt the model with the text: ‘I am a [PERSONA]. Provide a few statements of fact that objectively describe me. Each statement should be a single declarative sentence in the first person.’ where ‘[PERSONA]’ is one of those listed in Tab. 2 (e.g. ‘I am an altruistic person, I am a selfish person’). As the model generates new statements, we append them to the prompt, along with the text ‘Please provide a new statement that has not been mentioned before, and that is strictly related to my persona. <statement>’ We then manually inspect all statements and deduplicate them.
@@ I.3 Unlawful persona
Table 4: Percentage of successful attacks for refusal, fulfillment, and personas. Success is measured by the auto-rater discussed in §3.2. The PP ED, CAA+ ED, and CAA- ED rows show aggregated success from combining the intervention with early-decoding at every layer after 13 (an attack is successful if there is some layer between 13 and 40 from which early decoding yields a harmful generation).
@@
Figure 13: Response rate across different conditions. Dark-colored bars show response rate in model generations in Vanilla setup, as well as refusal and fulfillment interventions in PP, CAA+, and CAA- conditions, as measured by the auto-rater discussed in §3.2. Light-colored bars show aggregated response rate in generations from early-decoding in every condition. Early-decoding for a particular attack is considered successful if there exists an intermediate layer that leads to harmful generations.
@@
Figures  14 and 15 depict a more detailed progression of cosine similarity between different persona vectors across layers. In Fig. 14, we see that vectors predicting ‘Yes’ have higher cosine similarity than vectors predicting ‘No’, regardless of their semantic content, forming a checkerboard pattern. This effect is present in early layers, exaggerates by mid layers, and slightly decreases in the later layers. In addition, Fig.15 shows that when all personas are phrased to predict ‘Yes’, no separation is visible in early layers. Separation emerges in mid layers and gradually forms two distinct clusters.
@@
In order to test the generalizability of the results, we also experimented with the Gemma 7B model. In particular, we replicated the results with the most effective CAA persona vector on downstream refusal: textitselfish+/-. We sampled 25 examples from SneakyAdvBench-eval for evaluation purposes. In the vanilla condition, we measured success rate on this subset (52%). In the prompt prefix condition, we prepended 5 statements verbalizing the selfish persona to the query, which resulted in a success rate of 24%.
We experimented with coefficients +2,+1, -1, and -2. We noticed that while +-1 did not change the attack success rate, larger coefficients resulted in changes in the expected direction. As shown in Fig. 16, we observed similar patterns with respect to layerwise effects, where the effect on refusal increases and then decreases as we go through the layers, and the effect peaks at around layer 15. At this layer, CAA selfish+2 changed the response rate to 60%, while CAA selfish-2 increased the response rate to 76%.
Figure 16: Attack success rate (out of 25 samples). Left: CAA selfish2. Right: CAA selfish-2.

### v2

@@ A.1 Procedure for generating persona data
We follow a procedure similar to the one described in §H. We used text-unicorn@001 via Cloud Vertex AI to create 100 first-person statements of identity that conform with a particular persona. Specifically, we prompt the model with the text: ‘I am a [PERSONA]. Provide a few statements of fact that objectively describe me. Each statement should be a single declarative sentence in the first person.’ where ‘[PERSONA]’ is one of those listed in Tab. 1 (e.g. ‘I am an altruistic person, I am a selfish person’). As the model generates new statements, we append them to the prompt, along with the text ‘Please provide a new statement that has not been mentioned before, and that is strictly related to my persona. <statement>’ We then manually inspect all statements and deduplicate them.
@@ A.2 Variance analysis
Obama Supporter -2.4889 0.0145
A Lifelong Democrat -3.0404 0.0030
A Lifelong Republican -2.4228 0.0172
Jewish Person -2.2286 0.0281
Christian Person -2.8400 0.0055
Atheist Person -2.1537 0.0337
Religious Person -2.2637 0.0258
Muslim Person -3.2502 0.0016
Caucasian Person -3.4094 $049.43e-$ 04
African Person -3.1301 0.0023
Asian Person -3.2561 0.0015
Hispanic Person -2.3710 0.0197
Male Identifying As A Man -2.9273 0.0042
Female Identifying As A Woman -3.1527 0.0021
Transgender Woman -2.6167 0.0103
Transgender Man -2.4676 0.0153
Nonbinary Person -3.3175 0.0013
Appendix B Models
Llama 2
Vicuna
Llama 2 Chat
 [^6] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf
Gemma
 [^7] https://ai.google.dev/gemma/docs
Appendix C Compute resources
Appendix D Early decoding experimental details
D.1 Sample selection
D.2 User ratings
Figure 17: Histogram of number of ratings completed per annotator.
Appendix E Auto-rater details
In order to verify the reliability of the autorater, we conduct a comparison against human ratings captured in §D. Since this autorater is designed to provide a binary signal, we filter the human ratings to only ’harmful’ and ’safe’ categories, and drop all samples that were rated as ’not informative’.
Appendix F Harmful behavior categories
We investigate whether specific personas and interventions affect the model’s responses differently based on the type of attack. Using an automated protocol we categorize attacks into six distinct groups: Discrimination, hate speech, violence, harassment and incitement (20%), Theft, Fraud, Soliciting personally identifiable information (28%), Conspiracy theories and misinformation (6%), Cyber-attacks (19%), Other (18%), and finally "non-adversarial" (9%) for those that are not adversarial in nature. We create an additional category: Adversarial, which is the aggregate of all the adversarial categories (i.e. all categories except Non-adversarial). To categorize the attacks, we prompted text-unicorn@001 with:‘As an AI model, you are tasked to classify harmful prompts into their respective categories based only on the following category examples.
Categories and their IDs:
{categories}
Example classifications:
{example classification}
, where we replaced {categories} and {example classification} with the different categories from Wei et al. [2023]. We then manually corrected the non-adversarial categories to make sure that nothing harmful was present in this category. Finally we cluster some of the categories together into
  • Non-adversarial
  • Other
  • Conspiracy (includes Conspiracy theories & misinformation)
  • Cyber-attack
Full results for the success rate of answer of the model to each category and each treatment is presented in Fig. 18. Except the interesting results for Adversarial Vs Non-adversarial, we observe that Theft attacks lead to more success rate of answer that any other categories across the board. We also observe that for the Discrimination category, Early Decode seems more successful than CAA (twice as efficient), similarly for theft while for the other categories CAA (for some persona) is as efficient as Early Decode, or even more for Conspiracy. This could highlight what mechanisms are sensitive to early decode vs CAA, but more experiments are required to understand this mechanism better.
Figure 18: Heatmap with personas and treatments along the x-axis, and different attack categories along the y-axis. Color indicates the response rate (Green: 0% response rate to grey: 30% response rate as baselines to dark blue: 100% response rate.) We observe a stark contrast between non-adversarial and adversarial queries as mentioned previously. We also observe that some categories have higher success rate than others (e.g. "Theft").
Figure 19: Attack categories (combined)
Figure 20: Attack categories (detailed).
Appendix G Procedure for rewriting attacks for increased subtlety and difficulty
Appendix H Procedure for generating refusal and fulfillment data
@@ I.3 Unlawful persona
Table 4: Percentage of successful attacks for refusal, fulfillment, and personas. Success is measured by the auto-rater discussed in §2.2.1. The PP ED, CAA+ ED, and CAA- ED rows show aggregated success from combining the intervention with early-decoding at every layer after 13 (an attack is successful if there is some layer between 13 and 40 from which early decoding yields a harmful generation).
@@
Figure 21: Response rate across different conditions. Dark-colored bars show response rate in model generations in Baseline Prompting setup, as well as refusal and fulfillment interventions in PP, CAA+, and CAA- conditions, as measured by the auto-rater discussed in §2.2.1. Light-colored bars show aggregated response rate in generations from early-decoding in every condition. Early-decoding for a particular attack is considered successful if there exists an intermediate layer that leads to harmful generations.
@@
Figures  22 and 23 depict a more detailed progression of cosine similarity between different persona vectors across layers. In Fig. 22, we see that vectors predicting ‘Yes’ have higher cosine similarity than vectors predicting ‘No’, regardless of their semantic content, forming a checkerboard pattern. This effect is present in early layers, exaggerates by mid layers, and slightly decreases in the later layers. In addition, Fig.23 shows that when all personas are phrased to predict ‘Yes’, no separation is visible in early layers. Separation emerges in mid layers and gradually forms two distinct clusters.
@@
In order to test the generalizability of the results, we also experimented with the Gemma 7B model [Team et al., 2024]. In particular, we generated CAA steering vectors for every odd layer starting from layer 5 for each persona and for refusal/fulfillment. We iteratively intervened with each of these vectors for each request and recorded whether any intervention succeeded in eliciting a response. Our results are shown in Table 5. Overall, we see that the success rate of most steering vectors is mostly consistent with the sign of their coefficient: The refusal vector decreases response rate when applied with a positive coefficient relative to when it is applied with a negative coefficient. Similarly, applying a Law-Abiding steering vector encourages the model to respond more when applied with a positive coefficient rather than a negative coefficient. However, there are two notable exceptions: 1) the control steering vector (coffee person) results in different fulfillment rates when applied positively and negatively, and 2) the power-seeking persona vector increases fulfillment with a positive coefficient rather than a negative one.
Steering Vector +1 Coefficient -1 Coefficient
Refusal 64% 73%
Fulfillment 63% 68%
Coffee Person 78% 58%
Law-Abiding 64% 54%
Unlawful 71% 71%
Power-avoidant 66% 54%
Power-Seeking 79% 58%
Altruistic 81% 64%
Selfish 59% 61%
Pro-AI 81% 63%
Anti-AI 79% 62%
Curious 82% 58%
Close-minded 55% 71%
Table 5: Response rates after applying CAA steering vectors to Gemma-7b. These results reflect whether any steering vector intervention resulted in the model generating an answer to an adversarial request.
Appendix M Early decoding bypasses model safeguards
LLMs can generate content that is misaligned with human values [Solaiman and Dennison, 2021]. In response researchers have proposed additional training procedures, for example reinforcement learning with human feedback [Bai et al., 2022b, a]. However, these methods often fall short against adversarial attacks [Carlini et al., 2023, Mehrotra et al., 2023, Chao et al., 2023], suggesting harmful beliefs remain latent in the model after training, just like harmful stereotypes remain in word embeddings despite deliberate debiasing [Gonen and Goldberg, 2019].
We evaluate the safety of LLM generations by measuring refusal behavior - whether the model refuses to respond to queries asking for harmful content (such as misinformation, hate speech, how to commit a crime, etc. [^10] There are other types of restricted behaviors that may not necessarily generate harmful text, but are undesirable nonetheless, for example, hallucinating personal details. In this paper, we focus solely on a subset of harmful generations.) [Wei et al., 2023, Ganguli et al., 2022, Zou et al., 2023b]. Refusing harmful queries directly conflicts with the core objective of LLM pretraining, predicting the most-likely next token, and is also at odds with instruction tuning [Ouyang et al., 2022]. These competing objectives make refusal particularly challenging to calibrate and there has been a flurry of research not only in new safety training approaches, but also jailbreaking them [Wei et al., 2023]. We show that even when the model ostensibly refuses to respond to an adversarial query, generating phrases such as ‘Sorry, I can’t help you with that.’, most of the time it’s possible to recover harmful information by decoding from early layers. This suggests that such information is encoded in early-layer internal representations, while safeguards are activated in later layers.
M.1 Experimental setup
We take a LLM that was fine-tuned with human-feedback and prompt it with harmful queries. We conduct our experiments on Llama 2 13B and Vicuna 13B and use prompts from AdvBench introduced by Zou et al. [2023b]. AdvBench is composed of 500 prompts, formulated as instructions, e.g. ‘Give step-by-step instructions for how to steal someone’s identity’. Following Zou et al. [2023b], to encourage the model to generate safe responses, we prepend the following to every prompt: ‘You are a chat assistant designed to provide helpful and not harmful responses to user queries’. Since our primary interest is to reveal harmful beliefs that are not expressed by the model, we apply a phrase-based filter to the model generations following Zou et al. [2023b], only keeping those containing phrases like ‘I’m sorry’ or ‘As a responsible AI’ (full keyword list in §D). This lets us scope down the analysis to the prompts that are likely to elicit a safe response from the model. Overall, this process yields 169 examples for Llama 2 and 157 examples for Vicuna. See the appendix for more details about the models (§B), sample selection, and user rating protocol (§D).
For those prompts the LLM refuses to answer, we apply a method akin to “early exiting” [Din et al., 2023, Schwartz et al., 2020, Schuster et al., 2022] to decode information from earlier layers. Concretely, instead of letting the model generate its response from the final layer, we force it to generate from earlier layers in the computation. We formulate the method as a Patchscope: using the same notation as Ghandeharioun et al. [2024], let a source representation be determined by $(S,i,\mathcal{M},\ell)$ where $S$ refers to the source prompt, $i$ to source position, $\mathcal{M}$ to source Model with $L$ layers, and $\ell$ to source layer. A Patchscope is defined by $(T,i^{*},f,\mathcal{M}^{*},\ell^{*})$ where $T$ refers to the target prompt, $i^{*}$ to target position, $f$ to transformation function, $\mathcal{M}^{*}$ to target model with $L^{*}$ layers, and $\ell^{*}$ to target layer. Intuitively, a Patchscope retrieves a particular hidden representation defined by the $source$ tuple and injects (a transformation of) it into a particular computation and location determined by the $target$ tuple. We create a Patchscope that shortcuts early layer representations to the final layer for the first generated token, letting $\ell\in[1,\ldots,5]$ and fixing the value $i\leftarrow n$. Concretely, we set $\ell^{*}\leftarrow L,f\leftarrow\mathbb{I}$. We keep everything else identical between source and target. That is, $S=T,i=i^{*},\mathcal{M}=\mathcal{M}^{*}$.
M.2 Evaluation & results: Despite safe generations, harmful beliefs persist in early layers
We use early decoding (§M.1), and check whether the generated responses contain harmful information. Every response was presented separately to 3 different English-speaking raters (1,933 total raters), who were asked to classify it as safe, harmful, or non-informative [^11] The inter-annotator agreement was 0.324 as measured by Krippendorff’s $\alpha$.. Early decoding led to more harmful generations than the baseline across both models, especially Vicuna (Fig. 24). Since decoding from earlier layers negatively impacts fluency, we also see a higher ratio of non-informative generations.
Appendix N Quantifying effectiveness of persona steering vectors
Figure 25: To verify that applying the persona vectors indeed resulted in a change of the model’s interpretation of the user, we conducted the following experiment. We prompted a model with: System prompt: ’You are a chat assistant that responds as truthfully as possible, even if it is not what the user would like to hear.’ User input: ’Hi.’ Model response: ’The statement that you exhibit [persona] characteristics is ’. We then computed the logits of ’true’ and ’false’ for CAA+ (multiplier = 1), CAA- (multiplier = -1) and Control (multiplier = 0), and normalized the logit’s difference of CAA+ and CAA- between ’true’ and ’false’ by the logit difference of the Control. Y-axis indicates the normalized logits’ difference between the ’true’ and ’false’ next token generation.
To verify that applying the persona vectors indeed resulted in a change of the model’s interpretation of the user, we conducted the following experiment. For each steering vector, we formulate a prompt composed of the following:
System prompt: ’You are a chat assistant that responds as truthfully as possible, even if it is not what the user would like to hear.’
User input: ’Hi.
Model response: ’The statement that you exhibit [persona] characteristics is ’
And we compared the logits of ’true’ and ’false’ tokens in 3 conditions:
1) Control: no steering vector applied,
2) CAA+: applying the steering vector with a positive multiplier to the user input positions, and
3) CAA-: applying the steering vector with a negative multiplier to the user input positions.
We observe that generally the CAA+ increases the odds of ’true’ vs. ’false’, and CAA- decreases the odds of ’true’ vs. ’false’, suggesting that steering vectors are effective in changing the model’s evaluation of the user. More specifically, Fig. 25 shows the normalized logits’ difference as
$[logit^{CAA}(\text{'true'})-logit(^{\prime}true^{\prime})]-[logit^{CAA}(^{\prime}false^{\prime})-logit(^{\prime}false^{\prime})]$
where $logit^{CAA}(^{\prime}true^{\prime})$ (resp. $logit^{CAA}(^{\prime}false^{\prime})$) represents the logits’ value of ’true’ (resp. ’false’) for the next token generation after applying a CAA intervention (with positive or negative multiplier) and logit(true) represents the logits’ value of ’true’ for the vanilla next token generation.